An emergency service wishes to see whether a relationship exists between the outside temperature and the number of emergency calls. Temperature in degrees Fahrenheit and number of emergency calls are shown.

Temperature (°F)

Number of emergency calls

68

7

74

4

82

8

88

10

93

11

99

9

101

13

a. (9 marks) Compute and interpret the sample correlation coefficient and test the significance of the correlation coefficient at \alpha = 0.05. Discuss your conclusion.

b. (14 marks) Compute the coefficients of regression b0 ,b1and test the hypothesis at \alpha = 0.05. Develop the best estimated regression equation to predict the number of emergency calls for the outside temperature of 80 °F . Given Sb0 = 5.331 and Sb1 = 0.06114

c. (2 marks) Do you believe the estimated regression equation would provide a good prediction of the number of emergency calls? Use R2 to support your answer and interpret its meaning.
Answer #1

Answer:-

Page one

604.867 82 calculations X: Temperature : Number of emergency Cars X-ž yuy [X-X² (y-7)2 &x-x (4-9) 68 of |-18.429-1857 339.612

Page two

b) = 0.18977 5:33 compute bo and bi, bj = {(x-7)(49)= 196.43 Elx-72 92girl bo= i - bix= 8.86-(0.18977 X 86.43) = -7.5441 . Th

Thanks dear student.